OpenStudy (anonymous):

The product of (a - b)(a - b) is a perfect square trinomial. Sometimes Always Never

OpenStudy (anonymous):

Always. \((a-b)(a-b)=(a-b)^2=a^2-2ab+b^2\)

OpenStudy (anonymous):

Thanks!!